After 2 minutes on a treadmill, Jenny has a heart rate of 82. After 3 minutes she has a heart rate of 86. Assume that there is a linear equation that gives her heart rate h in terms of time on the treadmill t. Find the equation and use it to predict her heart tate after 10 minutes on the treadmill.
Answer by venugopalramana(3286)

After 2 minutes( T SAY ) on a treadmill, Jenny has a heart rate (H SAY )of 82. After 3 minutes she has a heart rate of 86. Assume that there is a linear equation that gives her heart rate h in terms of time on the treadmill t. Find the equation and use it to predict her heart tate after 10 minutes on the treadmill.
LET H=AT+B
WE HAVE FOR T=2,H=82...SO
82=2A+B.................................I
FOR T=3,H=86...SO
86=3A+B...............................II
EQN.II-EQN.I...
A=4...SUBSTITUTING IN EQN.I..
2*4+B=82
B=74
SO EQN.IS
H=4A+74
FOR T=10
H=4*10+74=114